# AdaptNC: Adaptive Nonconformity Scores for Uncertainty-Aware Autonomous Systems in Dynamic Environments

### v1

@@ C.3 Score Function Stability
The procedure described in Algorithm 1 creates a non-conformity score through a three step procedure: (1) Reweighting, (2) High-Density Region Identification, and (3) Convex Hull creation.
The reweighting procedure uses the expert weights $\gamma_{i}$ and the weights $w_{t}^{i}$ to create an exponentially weighted distribution $\mathcal{H}^{w}_{t}$. While we ordinarily have no guarantees on the evolution of these weights, under the assumption that $\eta_{t},\sigma_{t}\to 0$, the updates in the DtACI algorithm yield identical weights across timesteps $w_{t+1}^{i}=w_{t}^{i}$. Thus, the reweighted distributions $\mathcal{H}^{w}_{t}$ and $\mathcal{H}^{w}_{t+1}$ will be the same as $t\to\infty$.
For each distribution $\mathcal{H}^{w}_{t}$, the High-Density Region identification procedure uses a kernel density estimator to estimate the density, and then identifies the High-Density Region $\hat{R}_{t}$ that contains $1-\alpha$ mass of the estimated density. By Theorem B.1, the High-Density Region $\hat{R}_{t}$ will approximate the true high-density region $R_{\alpha}$ with high-accuracy as the number of samples $N$ and Monte Carlo iterations $M$ goes to $\infty$. Similarly, for $\mathcal{H}^{w}_{t+1}$, we obtain a High-Density Region $\hat{R}_{t+1}$, which will also approximate the true high-density region $R_{\alpha}$ with high-accuracy. The high density regions $\hat{R}_{t}=\hat{R}_{t+1}=R_{\alpha}$ will be equivalent as $N,M\to\infty$.
The definition of the Convex Hull implies uniqueness. Because of this, the Convex Hulls created from $\hat{R}_{t}$ and $\hat{R}_{t+1}$ will be the same as $t\to\infty$. Thus, the non-conformity score functions $s(\cdot;\theta_{t})$ and $s(\cdot;\theta_{t+1})$ will be the same as $t\to\infty$. ∎
Appendix D Experiments
D.1 Gaussian Mixture Model
This setting utilizes a Gaussian Mixture Model (GMM) with changing weights to provide intuition into the impact that changes in the non-conformity score have on the ideal miscoverage rate $\alpha^{*}_{t}$ in settings with distribution shift. The setting involves a mixture of two normal distributions, $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, mixed at time $t$ with a weight $w_{t}$. The random variable $z_{t}$ has the pdf $p(z_{t})=(1-w_{t})p(z_{t}|\mathcal{N}_{1})+w_{t}p(z_{t}|\mathcal{N}_{2})$. The data can be seen in Fig. 6. The mean and covariance of the normal distributions are:
@@ D.1 Gaussian Mixture Model
Figure 6: The 7000 points sampled from the Gaussian Mixture Model (GMM) data stream.
Figure 7: This figure shows the score functions evaluated as the weight parameter changes over time. We can see that the transition produces changes in the non-conformity scores.
We use the negative log-likelihood as the non-conformity score. For clarity, we define the non-conformity scores $s_{1}(z)=\textrm{NLL}(z|\mathcal{N}_{1})$, and $s_{2}(z|\mathcal{N}_{2})$. The empirical CDF $\hat{F}_{t,\mathcal{N}_{1}}(s)=\frac{1}{t}\sum_{\tau=1}^{t}s_{1}(z_{t})$, where $z_{t}$ is the observation at time $t$. The optimal quantile is the value $q^{*,\mathcal{N}_{1}}_{1-\alpha,t}=\inf_{q}\{q:\Pr[s_{1}(z_{t})\leq q]\geq 1-\alpha\}$, which induces the optimal miscoverage rate $\alpha^{*,\mathcal{N}_{1}}_{t}=\hat{F}^{-1}_{t,\mathcal{N}_{1}}(q^{*,\mathcal{N}_{1}}_{1-\alpha,t})$. These quantities are similarly defined for the non-conformity score $s_{2}$. Note that the optimal quantile and the optimal miscoverage rate are only dependent on the two score functions, and do not assume that the data are drawn from these distributions. We can see the value of the non-conformity scores as the weight parameter changes in Fig. 7. Based on the non-conformity scores, we can also calculate the values of our optimal $\alpha^{*}_{t}$ under each score function.
Figure 8: This figure shows the impact of changing non-conformity scores under a shifting GMM. Top: We compute the induced theoretical $(1-\alpha)$ quantiles for each score function, which shift in opposite directions as the dominant mixture component changes. Middle: The corresponding optimal miscoverage rates $\alpha_{t}^{\star}$ implied by each score function, illustrating that distribution shift can make the ideal miscoverage rate deviate substantially from the nominal $\alpha=0.1$ (red dashed), purely due to the change in score distribution. Bottom: The difference between the two $\alpha_{t}^{*}$.
Figure 9: This figure shows the local and global coverage performance under a shifting GMM. We compare DtACI using the fixed score $s_{1}$, DtACI using the fixed score $s_{2}$, and AdaptNC. Top: Local (sliding-window) coverage with window size $w=200$ against the $1-\alpha=0.90$ target, highlighting that a mismatched score can lead to sharp under-coverage during the transition while AdaptNC remains near-target. Middle: Global coverage over time, showing persistent bias for DtACI when the score is misaligned with the dominant mixture component. Bottom: The mixture weight $w_{t}$ on $\mathcal{N}_{2}$, which induces the distribution shift driving the coverage changes.
This is shown in Fig. 9. The figure depicts the difference in $\alpha^{*}$ over time. This quantity is of interest to us, as the theoretical bound Remark 5.2 is dependent on the difference between subsequent $\alpha^{*}_{t}$. While the change in $\alpha^{*,\mathcal{N}_{1}}_{t}$ and $\alpha^{*,\mathcal{N}_{2}}_{t}$ is smooth, the problem we encounter in settings with score function change is that the difference between the two $\alpha^{*}_{t}$ can be large. This indicates that if we were to switch score functions at the time of distribution shift, we would incur a large regret penalty, as our miscoverage rate would be poorly calibrated to the new score function. This motivates the need for adaptive miscoverage rates in settings with distribution shift and changing score functions. The counterfactual replay mechanism we propose in Section 4 is designed to mitigate this issue by replacing a potentially abrupt change in $\alpha^{*}_{t}$ due to a score function shift, to the local evolution of $\alpha^{*,\mathcal{N}_{2}}_{t}$, which is much smoother.
This is evidenced in Fig. 9, where we compare the local and global coverage of DtACI using fixed score functions $s_{1}$ and $s_{2}$, against AdaptNC which adapts to the changing score function. We can see that when the score function is misaligned with the dominant mixture component, DtACI suffers from sharp under-coverage during the transition period, as well as persistent bias in global coverage. In contrast, AdaptNC is able to maintain near-target local and global coverage throughout the data stream, demonstrating the effectiveness of our adaptive miscoverage rate approach in handling distribution shift and changing non-conformity scores.
D.2 Indoor Localization

### v2

@@ D.2 HDR and Score-Stability Proofs
By the assumed stability of the weighted residual distributions and the assumed relation between $D_{\phi}$ and the residual density class,
$\|f_{t}-f_{\star}\|_{\infty}\to 0.$
By uniform KDE consistency,
$\|\widehat{f}_{t}-f_{t}\|_{\infty}\xrightarrow{p}0.$
Therefore,
$\|\widehat{f}_{t}-f_{\star}\|_{\infty}\xrightarrow{p}0.$
The Monte Carlo threshold $\widehat{\tau}_{t}$ is the empirical quantile of the density scores $\widehat{f}_{t}(Z)$, where $Z$ is sampled from the KDE. Since $M_{t}\to\infty$, the empirical density-score CDF converges uniformly to its population counterpart, by the Glivenko–Cantelli theorem. Together with the regularity of the density-score quantile and the convergence $\widehat{f}_{t}\to f_{\star}$, this gives
$|\widehat{\tau}_{t}-\tau_{\star}|\xrightarrow{p}0.$
Define
$\varepsilon_{t}:=\|\widehat{f}_{t}-f_{\star}\|_{\infty}+|\widehat{\tau}_{t}-\tau_{\star}|.$
Then $\varepsilon_{t}\xrightarrow{p}0$.
If $z\in\widehat{R}_{t}$, then $\widehat{f}_{t}(z)\geq\widehat{\tau}_{t}$, and hence
$f_{\star}(z)\geq\widehat{f}_{t}(z)-\|\widehat{f}_{t}-f_{\star}\|_{\infty}\geq\widehat{\tau}_{t}-\|\widehat{f}_{t}-f_{\star}\|_{\infty}\geq\tau_{\star}-\varepsilon_{t}.$
Therefore,
$\widehat{R}_{t}\subseteq\{z\in\Omega:f_{\star}(z)\geq\tau_{\star}-\varepsilon_{t}\}.$
Similarly, if $z\in R_{\star}$, then $f_{\star}(z)\geq\tau_{\star}$. Points in $R_{\star}$ with $f_{\star}(z)\geq\tau_{\star}+\varepsilon_{t}$ satisfy
$\widehat{f}_{t}(z)\geq f_{\star}(z)-\|\widehat{f}_{t}-f_{\star}\|_{\infty}\geq\tau_{\star}+\varepsilon_{t}-\|\widehat{f}_{t}-f_{\star}\|_{\infty}\geq\widehat{\tau}_{t},$
and therefore belong to $\widehat{R}_{t}$. Thus, any point of $R_{\star}$ that may fail to belong to $\widehat{R}_{t}$ must lie in the shrinking level band
$\{z\in\Omega:|f_{\star}(z)-\tau_{\star}|\leq\varepsilon_{t}\}.$
By the regular-boundary assumption, these level bands shrink to $\partial R_{\star}$ in Hausdorff distance. More precisely, for all sufficiently small $\varepsilon$, every point satisfying
$|f_{\star}(z)-\tau_{\star}|\leq\varepsilon$
lies within distance $O(\varepsilon)$ of the boundary $\{z:f_{\star}(z)=\tau_{\star}\}$. Consequently,
$d_{H}(\widehat{R}_{t},R_{\star})\leq C\varepsilon_{t}$
for some constant $C>0$ and all sufficiently large $t$, with probability tending to one. Since $\varepsilon_{t}\xrightarrow{p}0$, we conclude that
$d_{H}(\widehat{R}_{t},R_{\star})\xrightarrow{p}0.$
Lemma D.6 (Convex-hull stability).
Let $A_{m},A_{\star}\subset\mathbb{R}^{p}$ be nonempty compact sets.
$d_{H}(A_{m},A_{\star})\to 0\implies d_{H}(\operatorname{conv}(A_{m}),\operatorname{conv}(A_{\star}))\to 0.$
Proof.
Let $\rho_{m}=d_{H}(A_{m},A_{\star})$. For any $x\in\operatorname{conv}(A_{m})$, write
$x=\sum_{\ell=1}^{L}\lambda_{\ell}a_{\ell},\qquad a_{\ell}\in A_{m},\quad\lambda_{\ell}\geq 0,\quad\sum_{\ell=1}^{L}\lambda_{\ell}=1.$
For each $a_{\ell}$, choose $a_{\ell}^{\star}\in A_{\star}$ such that $\|a_{\ell}-a_{\ell}^{\star}\|_{2}\leq\rho_{m}$. Then
$x^{\star}:=\sum_{\ell=1}^{L}\lambda_{\ell}a_{\ell}^{\star}\in\operatorname{conv}(A_{\star}),$
and
$\|x-x^{\star}\|_{2}\leq\sum_{\ell=1}^{L}\lambda_{\ell}\|a_{\ell}-a_{\ell}^{\star}\|_{2}\leq\rho_{m}.$
where the last inequality follows from the definition of the Hausdorff distance. Thus every point of $\operatorname{conv}(A_{m})$ lies within $\rho_{m}$ of $\operatorname{conv}(A_{\star})$. The reverse inclusion is identical. Hence
$d_{H}(\operatorname{conv}(A_{m}),\operatorname{conv}(A_{\star}))\leq\rho_{m}.$
Corollary D.7.
In the setting of Lemma˜D.6,
$d_{H}(\operatorname{conv}(A_{m}),\operatorname{conv}(A_{\star}))\leq d_{H}(A_{m},A_{\star}).$
Because halfspace representations of a convex set are not unique, we state score stability for the canonical support-function score induced by the fitted convex template. Equivalently, the implemented halfspace score should be normalized to this canonical representation before comparing scores across updates. Our implementation, which relies on QuickHull [2], performs this normalization.
Lemma D.8 (Score stability from hull stability).
For a compact convex set $K\subset\mathbb{R}^{p}$, define
$s_{K}(x,y):=\sup_{\|v\|_{2}=1}\{v^{\top}(y-h(x))-H_{K}(v)\},\qquad H_{K}(v):=\sup_{z\in K}v^{\top}z.$
If
$d_{H}(K_{t},K_{\star})\to 0,$
then
$\sup_{(x,y)\in\mathcal{A}}|s_{K_{t}}(x,y)-s_{K_{\star}}(x,y)|\to 0$
for a set $\mathcal{A}\subseteq\mathcal{X}\times\mathcal{Y}$.
Proof.
For any residual z=yh(x)z=y-h(x),
$|s_{K_{m}}(x,y)-s_{K_{\star}}(x,y)|$ $=\left|\sup_{\|v\|_{2}=1}\{v^{\top}z-h_{K_{m}}(v)\}-\sup_{\|v\|_{2}=1}\{v^{\top}z-h_{K_{\star}}(v)\}\right|$
$\leq\sup_{\|v\|_{2}=1}|H_{K_{m}}(v)-H_{K_{\star}}(v)|.$
For compact convex sets,
$\sup_{\|v\|_{2}=1}|H_{K_{m}}(v)-H_{K_{\star}}(v)|=d_{H}(K_{m},K_{\star}).$
Therefore,
$|s_{K_{m}}(x,y)-s_{K_{\star}}(x,y)|\leq d_{H}(K_{m},K_{\star}).$
Taking the supremum over $(x,y)\in\mathcal{A}$ gives the result. ∎
See 5.2
Proof.
By Lemma D.5,
$d_{H}(\widehat{R}_{m},R_{\star})\xrightarrow{p}0.$
By Lemma D.6,
$d_{H}(K_{m},K_{\star})=d_{H}(\operatorname{conv}(\widehat{R}_{m}),\operatorname{conv}(R_{\star}))\xrightarrow{p}0.$
By Lemma D.8,
$\sup_{(x,y)\in\mathcal{A}}|s_{K_{m}}(x,y)-s_{K_{\star}}(x,y)|\xrightarrow{p}0.$
Finally, by the triangle inequality,
$\sup_{(x,y)\in\mathcal{A}}|s_{K_{m+1}}(x,y)-s_{K_{m}}(x,y)|$
$\qquad\leq\sup_{(x,y)\in\mathcal{A}}|s_{K_{m+1}}(x,y)-s_{K_{\star}}(x,y)|+\sup_{(x,y)\in\mathcal{A}}|s_{K_{m}}(x,y)-s_{K_{\star}}(x,y)|.$
Both terms converge to zero in probability, so the successive score functions are asymptotically stable. ∎
Proof of Corollary˜5.3.
Suppose
$\sup_{(x,y)\in\mathcal{A}}|s_{t}(x,y)-s_{\star}(x,y)|\leq\delta_{t}.$
Under Assumption D.4, for every fixed threshold $q$,
$\left|\operatorname{Vol}(C_{t}(x;q))-\operatorname{Vol}(C_{\star}(x;q))\right|\leq L_{\mathrm{vol}}\delta_{t}.$
Since $\delta_{t}\to 0$ in probability by Theorem˜5.2, the prediction-region volume stabilizes in probability. ∎
D.3 Long-Run Coverage under Replay
Proof of Theorem˜5.4.
Expanding the perturbed recursion gives
$\alpha_{T+1}^{i}=\alpha_{1}^{i}+\gamma_{i}\sum_{t=1}^{T}(\alpha-\operatorname{\textbf{err}}_{t}^{i})+\sum_{t=1}^{T}\Delta_{t}^{i}.$
Rearranging,
$\frac{1}{T}\sum_{t=1}^{T}\operatorname{\textbf{err}}_{t}^{i}-\alpha=\frac{\alpha_{1}^{i}-\alpha_{T+1}^{i}}{T\gamma_{i}}+\frac{1}{T\gamma_{i}}\sum_{t=1}^{T}\Delta_{t}^{i}.$
The first term vanishes because the expert states are bounded. The second term vanishes by the assumed sublinear cumulative perturbation.
$\left|\frac{1}{T\gamma_{i}}\sum_{t=1}^{T}\Delta_{t}^{i}\right|\leq\frac{1}{T\gamma_{i}}\sum_{t=1}^{T}|\Delta_{t}^{i}|\to 0.$
Hence:
$\frac{1}{T}\sum_{t=1}^{T}\operatorname{\textbf{err}}_{t}^{i}\to\alpha.$
Proof of Theorem˜5.5.
For the analysis, define the pre-update replay trajectory as the counterfactual trajectory obtained by replaying $\mathcal{R}_{m}$ using $s_{m}^{-}$, and the post-update replay trajectory as the trajectory obtained by replaying the same window using $s_{m}^{+}$, both initialized from the same expert state. Let
$e_{m,r}^{i,-}:=\mathbf{1}\{s_{m}^{-}(X_{r},Y_{r})>q_{m,r}^{i,-}\},\qquad e_{m,r}^{i,+}:=\mathbf{1}\{s_{m}^{+}(X_{r},Y_{r})>q_{m,r}^{i,+}\}.$
At non-update times, $\Delta_{t}^{i}=0$. At update time $t_{m}$, the difference between the replayed post-update expert state and the corresponding pre-update expert state is controlled by the number of replayed observations whose coverage indicators change:
$|\Delta_{t_{m}}^{i}|\leq\gamma_{i}\sum_{r\in\mathcal{R}_{m}}|e_{m,r}^{i,+}-e_{m,r}^{i,-}|.$
Now observe that the indicators can differ only if the replayed score lies near the pre-update threshold. Indeed, if
$|s_{m}^{-}(X_{r},Y_{r})-q_{m,r}^{i,-}|>(1+L_{q})\varepsilon_{m},$
then
$|s_{m}^{+}(X_{r},Y_{r})-s_{m}^{-}(X_{r},Y_{r})|\leq\varepsilon_{m}$
and
$|q_{m,r}^{i,+}-q_{m,r}^{i,-}|\leq L_{q}\varepsilon_{m}$
together imply that $s_{m}^{+}(X_{r},Y_{r})$ and $s_{m}^{-}(X_{r},Y_{r})$ remain on the same side of their respective thresholds. Hence
$|e_{m,r}^{i,+}-e_{m,r}^{i,-}|\leq\mathbf{1}\left\{|s_{m}^{-}(X_{r},Y_{r})-q_{m,r}^{i,-}|\leq(1+L_{q})\varepsilon_{m}\right\}.$
Therefore
$\frac{1}{T}\sum_{t=1}^{T}|\Delta_{t}^{i}|\leq\frac{\gamma_{i}}{T}\sum_{m:t_{m}\leq T}\sum_{r\in\mathcal{R}_{m}}\mathbf{1}\left\{|s_{m}^{-}(X_{r},Y_{r})-q_{m,r}^{i,-}|\leq(1+L_{q})\varepsilon_{m}\right\}.$
The assumed averaged boundary condition implies that the right-hand side converges to zero. The final coverage claim follows from the perturbed single-expert ACI guarantee. ∎
D.4 Coverage Shock from Score Updates
Proposition D.9.
Let $\mathcal{D}_{t}$ and $\mathcal{D}_{t+1}$ be the joint distributions of inputs and observations $(X_{t},Y_{t})$ at times $t$ and $t+1$. Assume that the total variation distance ($\mathbf{TV}$) between the two distributions is bounded $\mathbf{TV}(\mathcal{D}_{t},\mathcal{D}_{t+1})\leq\delta$. For a fixed NCSF $s(\cdot)$, and a fixed score parameter $\theta$ let $S_{\theta,t}$ be the random variable $s(X,Y,\theta)$ where $(X,Y)\sim\mathcal{D}_{t}$, and $S_{\theta,t+1}$ be the random variable where $(X,Y)\sim\mathcal{D}_{t+1}$ Then, the $\mathbf{TV}$ distance between the score distributions will be less than $\delta$: $\mathbf{TV}(S_{\theta,t},S_{\theta,t+1})\leq\delta$
Proof.
This follows from the Data Processing Inequality [21, Theorem 7.4] with NCSF $s(\cdot;\theta)$ as channel, and $f$-divergence as $\mathbf{TV}$ distance. ∎
Appendix E Experiments
E.1 Gaussian Mixture Model: A Minimal Example of Coverage Shock
Figure 5: Score-induced coverage shock in GMM example. The top panel shows the gap between history-relative effective miscoverage levels induced by the two diagnostic scores, while the bottom panel shows the rate of change of the mixture weight. Despite smooth distributional changes, altering the score can induce large shifts in the calibration state that DtACI must track.
The replay mechanism in AdaptNC is motivated by a simple failure mode: the state of an adaptive thresholding algorithm is tied to the nonconformity score used to map past residuals into ranks and coverage errors. When the score function changes, the same recent residuals can induce a different score distribution and therefore a different calibration state. Thus, even if the underlying residual distribution changes smoothly, a score update can produce a transient mismatch between the inherited threshold state and the updated score. We refer to this transient mismatch as coverage shock. The following Gaussian mixture example isolates this effect in a setting where the only source of nonstationarity is a smooth change in mixture weights.
Data-generating process.
At time $t$, a residual sample $z_{t}\in\mathbb{R}^{2}$ is drawn from a two-component Gaussian mixture
$p_{t}(z)=(1-w_{t})p_{1}(z)+w_{t}p_{2}(z),$ (11)
where $w_{t}$ varies smoothly over time. Thus the stream begins in a regime dominated by $\mathcal{N}_{1}$ and gradually transitions to a regime dominated by $\mathcal{N}_{2}$. The two components are
@@ E.1 Gaussian Mixture Model: A Minimal Example of Coverage Shock
The resulting stream is shown in Fig.˜6.
Figure 6: The 7000 samples drawn from the shifting Gaussian mixture stream. Color indicates time, showing the transition from the $\mathcal{N}_{1}$-dominated regime to the $\mathcal{N}_{2}$-dominated regime.
Figure 7: Negative log-likelihood scores under the two component scores as the mixture weight changes. The smooth mixture transition induces a systematic change in the score distribution.
Score-dependent reference levels.
We compare two fixed nonconformity scores, each matched to one component:
$s_{a}(z)=-\log p_{a}(z),\qquad a\in\{1,2\}.$ (12)
These scores are used as interpretable diagnostic endpoints: $s_{1}$ is matched to the early regime and $s_{2}$ is matched to the late regime. AdaptNC does not literally switch between these two scores; rather, the pair illustrates how a change in score geometry can change the calibration state even on the same residual stream.
For each score $s_{a}$, let
$q_{a,t}^{\mathrm{mix}}=\inf\left\{q:\Pr_{Z\sim p_{t}}\left[s_{a}(Z)\leq q\right]\geq 1-\alpha\right\}$ (13)
be the nominal $(1-\alpha)$ score quantile under the instantaneous mixture $p_{t}$. In the experiment, this quantity is estimated by Monte Carlo from the current mixture distribution. We then evaluate where this instantaneous quantile falls relative to the accumulated score history,
$\widehat{F}^{\mathrm{hist}}_{a,t}(u)=\frac{1}{t+1}\sum_{\tau=0}^{t}\mathds{1}\{s_{a}(z_{\tau})\leq u\},\qquad\widehat{\alpha}_{a,t}^{\star}=1-\widehat{F}^{\mathrm{hist}}_{a,t}\!\left(q_{a,t}^{\mathrm{mix}}\right).$ (14)
The quantity $\widehat{\alpha}_{a,t}^{\star}$ is a diagnostic effective miscoverage level: it is the miscoverage rate that would make the accumulated score history place its $(1-\widehat{\alpha}_{a,t}^{\star})$ quantile at the instantaneous mixture quantile. Because both $q_{a,t}^{\mathrm{mix}}$ and $\widehat{F}^{\mathrm{hist}}_{a,t}$ depend on the score map, changing from $s_{1}$ to $s_{2}$ can induce a large jump in $\widehat{\alpha}_{a,t}^{\star}$ even when $w_{t}$ changes smoothly.
Figure 8: Coverage shock in the shifting GMM. Top: The nominal $(1-\alpha)$ score quantiles for the two diagnostic scores evolve in opposite directions as the dominant mixture component changes. Middle: The corresponding effective miscoverage levels $\widehat{\alpha}_{a,t}^{\star}$ deviate from the nominal target $\alpha=0.1$ and from one another. Bottom: The gap between these levels quantifies the change in calibration state induced by changing scores.
Figure 9: Coverage behavior under the shifting GMM. The fixed $s_{1}$ baseline undercovers after the mixture shifts, while the fixed $s_{2}$ baseline is conservative. AdaptNC remains closer to the target than the under-covering fixed score and avoids the persistent conservatism of the late-regime score by adapting its score and replaying recent observations.
Why this motivates replay.
Fig.˜9 shows that a smooth change in mixture weight can still create a large separation between the calibration reference levels associated with different scores. If a score-adaptive method updates its score without recalibrating the threshold state, DtACI continues from a state shaped by old score observations while receiving future observations through a new score. In the regret bound of Remark˜D.2, this appears as a large $|\alpha_{t}^{\star}-\alpha_{t-1}^{\star}|$ term, i.e., coverage shock.
Replay mitigates this effect by reconstructing the recent DtACI state under the updated score. Instead of applying the new score to future observations while retaining a threshold state learned from the old score, AdaptNC rescores the recent window, recomputes the corresponding ranks and coverage errors, and reruns the threshold updates. In the GMM example, this replaces a counterfactual abrupt jump between the diagnostic endpoint scores with a recalibration step under the updated score used by AdaptNC.
This behavior is reflected in Fig.˜9. When a fixed score becomes mismatched with the dominant mixture component, DtACI either undercovers or becomes conservative. AdaptNC adapts the score and uses replay to realign the threshold state, yielding more stable near-target coverage through the transition. The example is intentionally simple: its purpose is not to model a robotics task, but to isolate the mechanism by which score changes create coverage shock and to motivate replay as the corresponding state-realignment step.
E.2 Indoor Localization